# mO Revanth[#1]

### v1

@@ 4.3 Multi-level Contrastive Learning
Given a dataset with $N$ samples, the $i^{th}$ sample is denoted as $T_{i}$ and assigned with three-level labels indicating its source.
  • $x$: if $T_{i}$ is fully AI-generated, $x_{i}=0$, otherwise $x_{i}=1$;
  • $y$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
  • $z$: an indicator of a specific LLM family.
@@
For AI-generated text ($x=0$), the similarity between $T_{i}$ and another AI-generated text $T_{j}$ is greater than that with a human-written or collaborative text $T_{k}$ ($x=1$):
@@
If $x=0$, then the text is fully AI-generated. For this case, we do not consider $y$ since AI-generated text is considered as non-AI-collaboration. We can imply that the similarity between two texts written by the same AI is higher than that of two different AI families. Hence:
@@
The reverse condition (with $y=1$) is also true. We can conclude that:
@@
For cases where all samples are human-AI collaborative texts ($x,y=1$), two texts with involvements of the same AI family tend to be more similar than the ones that involve contributions from different families. That is:
@@
$\begin{cases}\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=0,z_{i}=z_{j}\neq z_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j}\neq y_{k};\\ \begin{split}\sigma(i,j)>\sigma(i,k),\forall&x_{i}=1,y_{i,j,k}=1,\\ &z_{i}=z_{j},z_{i}\neq z_{k}\\ \end{split}\end{cases}$
To enforce the similarity constraints outlined in Eq (6), we build upon the SimCLR framework Chen et al. (2020) and introduce a strategy for defining both positive and negative sample pairs, which forms the basis of our contrastive learning loss. Departing from traditional contrastive losses that rely on a single positive sample, our approach considers a group of positive instances that satisfy specific criteria. The similarity between the anchor and positive samples is computed as the average similarity across this entire positive set. For negative samples, we follow the methodology used in SimCLR. The resulting contrastive loss, expressed in Eq (7), involves $q$ as the anchor sample, $K^{+}$ as the positive sample set, $K^{-}$ as the negative sample set, $\tau$ as the temperature parameter, and $N_{K^{+}}$ as the number of positive samples.
@@
Therefore, due to the last inequality in Eq (6) only specifying a case ($y=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$. Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$. We set $\gamma=\delta=1$, then $\zeta=2,\alpha=\beta=2$.
This approach encourages the model to capture subtle and detailed features from different sources. As a result, the model becomes more adept at recognizing variations in writing styles. This capability enhances its accuracy and strengthens its ability to generalize when detecting AI-generated text.
@@ 4.4 Multi-task Auxiliary Learning
Multi-task learning Caruana (1997) allows a model to learn several tasks concurrently by sharing relevant information across them. This joint learning process helps the model develop more general and distinctive features. As a result, it improves the model’s ability to generalize to new data. Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer. This classifier is responsible for binary classification, determining whether a given query text was written by a human or generated by a LLM. Denote that the probability of $i^{th}$ sample with label $x_{i}=0$ is $p_{i}$. To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, defined as follows:
@@
Figure 2: Inference architecture. (a) embed the input text into embedding vector using the fine-tuned encoder; (b) use Fuzzy KNN algorithm to cluster, retrieving which cluster the input text belongs to (see more in Appendix E). (c) The stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder. If the input text is unseen, we embed the unseen text and save into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
@@ 4.5 Handling Unseen Data without Retraining
Unseen data, whether from an unseen domain or generated by an unfamiliar model, remains a significant challenge even for state-of-the-art AI-generated text detection methods, due to the increasing capabilities of large language models. We tried to use the model classifier on its own, but we ended up by using a vector database along with Fuzzy k-Nearest Neighbors, which is illustrated in Figure 2. The results were written in Appendix E. Specifically, when dealing with unseen data, we use our trained model to embed these samples and add them to our existing vector database. Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without the need retraining.
@@ 5 Experiments
In this section, we first describe datasets and baselines used, followed by two experiments evaluating FAID in classifying a text between human, AI and human-AI; and in identifying specific generators.
@@
In-domain, Known generator FAIDSet LLM-DetectAIve 94.34 94.45 93.79 94.10 0.1888 0.1107
SVM (N-grams) 84.41 84.53 82.93 83.62 0.5630 0.2916
@@
Unseen domain, Unseen generator Unseen domain LLM-DetectAIve 52.83 47.31 64.62 53.28 0.4733 0.4722
SVM (N-grams) 39.28 43.13 29.41 22.77 0.7689 0.6611
@@
Unseen generators LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
SVM (N-grams) 74.19 61.88 50.17 55.17 0.4324 0.3162
@@
Unseen domain and Unseen generators LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
SVM (N-grams) 34.50 39.38 26.67 21.70 0.9186 0.7429
@@
Table 1: Performance with three labels. For in-domain and known generator datasets, we train detectors on training set and evaluate using test set. For unseen data evaluation, we use the detector trained on FAIDSet, and evaluate on the unseen datasets. The best results are in bold and the second best are underlined.
@@ 5.1 Datasets
In addition to FAIDSet, for in-domain evaluation, we used another two datasets:
LLM-DetectAIve Abassy et al. (2024) encompasses various domains including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews. The original human-written and machine-generated texts were augmented using a variety of LLMs (e.g., Llama 3, Mixtral, Gemma) to create 236,000-example dataset with two new labels: machine-written then machine-humanized, and human-written then machine-polished.
HART Bao et al. (2025) is a dataset spans domains of student essays, arXiv abstracts, story writing, and news articles. It has four categories: human-written texts, LLMs-refined texts, AI-generated texts, and humanized AI-generated texts. They further generated 21,500 cases based on this dataset for unbalanced labels.
To evaluate generalization ability of FAID on unseen scenarios, we collected the following data:
Unseen domain: We create a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written. These essays were then used to generate human-AI collaborative and AI-generated texts, using the same models with FAIDSet.
Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for remaining labels using the three new LLM families: Qwen 2, Mistral, and Gemma 2.
Unseen domain & generators: Based on the human-written IELTS essays above, instead of the original four LLMs, we used the same LLM families as Unseen generator test set to generate data for the AI and human-AI labels, following the same process as with the unseen domains.
@@ 5.2 Baselines
Abassy et al. (2024) fine-tuned a RoBERTa Liu et al. (2019a) on fine-grained English AI-generated texts with learning rate of $2\times 10^{-5}$ for 10 epochs.
Support Vector Machine with N-grams:
Prova (2024) utilized SVM with N-gram features for binary AI text detection (Human vs. AI). We adapted this approach to a three-class detection. We first generated word-level trigram features using Count Vectorizer, and then used a Linear SVM to train with the gold labels to guide the classification.
@@
Chen et al. (2023) utilized T5-Sentinel to classify text into two labels (Human vs. AI). This approach leverages the T5 model to predict the conditional probability of the next token, integrating the classification task into a sequence-to-sequence completion framework. This unique method directly leverages the strength of encoder-decoder model, which are trained to understand context and generate coherent text.
5.3 Human, AI or Human-AI?
Table 1 shows the performance of FAID and three baselines in four evaluation settings. FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domain, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii). Taking advantage of pretrained language models (i.e., RoBERTa, T5), they outperform SVM in most cases, particularly on HART dataset.
FAID’s approach enhances the generalization performance over unseen domains and generators, compared with other methods. From the absolute accuracy, we can find that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55% respectively. The result suggests that FAID is an effective method to address the multilingual fine-grained AI-generated text detection task, improving the performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to AI families, rather than overfitting to surface-level artifacts.
@@ 5.4 Identifying Different Generators
We further examine the detector to identify specific AI generators which were seen during training. The goal of FAID is not only to detect AI involvement but also to identify the specific AI model or family applied, treating them as distinct authors. As shown in Table 2, across the three datasets, FAID consistently achieves higher performance compared to other baselines in almost all metrics. LLM-DetectAIve achieves comparable results with our detector FAID on the test set of LLM-DetectAIve dataset, except that its precision is slightly lower. FAID’s high performance when dealing with text originating from diverse known generators within these datasets indicates that FAID learned unique writing patterns and features of different AI generators by leveraging multi-level contrastive learning.
6 Conclusion and Future Works
In this work, we collected a multilingual fine-grained AI-generated text detection dataset FAIDSet with $\sim$84,000 examples. We further introduced FAID, a fine-grained AI-generated text detection framework aiming to distinguish between human-written, AI-generated, and human-AI collaborative texts. It combines multi-level contrastive learning and multi-task auxiliary objectives, with the perspective of treating AI model families as stylistic authors. FAID captures subtle writing patterns that generalize effectively across domains and evolving AI systems, where Fuzzy KNN-based inference method and training-free incremental adaptation enhance the robustness, enabling strong performance on unseen data.
Experiments demonstrate that FAID consistently outperforms other strong baselines across multiple datasets and configurations. Its capability to recognize nuanced forms of collaborative texts and the ability to adapt to new generative models make it particularly suited for real-world applications, improving transparency in AI-assisted writing.
In the future, we plan to expand FAIDSet by incorporating more languages, generators and domains, especially low-resource languages and informal text genres such as social media content or student essays, to further enhance cross-lingual and domain robustness. We also continue to investigate to deal with AI-generated text or more complex human-AI collaborative texts (i.e., from two or more AI families, generated by multiple AI families then edited by human) to deal with human’s behaviors in using AI for writing.
@@ 7 Limitations
While FAID demonstrates strong performance and generalization across various domains and AI models, several limitations remain. First, although our dataset is multilingual and multi-domain, it is still limited in terms of low-resource languages and niche writing domains, which may affect performance in those contexts. Second, our framework is based on the fact that texts produced by AI models from the same family share similar stylistic features. However, this assumption may break down when a single text is influenced by multiple AI systems – such as when a human uses different models for drafting, rewriting, and polishing. In such cases, the resulting style may blend traits from multiple sources, making it more difficult to attribute authorship to a single AI family or clearly distinguish collaboration boundaries.
8 Ethics and Broader Impacts
@@
A primary ethical consideration is the data license. We reused existing dataset for our research: LLM-DetectAIve, HART and IELTS Writing, which have been publicly released under clear licensing agreements. We adhere to the intended usage of these dataset licenses.
@@
FAIDSet streamlines both the creation and rigorous testing of FAID. By spotting AI-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors. More broadly, it supports efforts to prevent the misuse of generative technologies in areas such as credential falsification. Detecting AI-crafted content across different languages can be tricky, due to language’s unique grammar and style. By enabling robust, multilingual and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly. At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and limits of generative AI.
@@
AI-generated text detection systems may inadvertently encode or amplify biases present in training data. FAIDSet has been carefully constructed to include diverse domains and languages to reduce such biases. Nonetheless, we encourage ongoing auditing and benchmarking of fairness across populations and writing styles, and welcome community feedback for further improvements.
@@
In this work, to promote open research and community contributions, we published our code and data to the public.

### v2

@@ 4.2 Multi-level Contrastive Learning
Given a dataset with $N$ samples, each sample is a text unit (paragraph/segment). The $i^{th}$ sample is denoted as $T_{i}$. Per sample, we assigned three-level labels indicating its source:
  • $x_{i}\in\{0,1\}$: if $T_{i}$ is fully LLM-generated, $x_{i}=0$, otherwise $x_{i}=1$;
  • $y_{i}\in\{0,1\}$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
  • $z_{i}$: an indicator of a specific LLM family.
@@
For LLM-generated text, the similarity between $T_{i}$ and another LLM-generated text $T_{j}$ is greater than that with a human-written or collaborative text $T_{k}$:
@@
If $x_{i}=0$, then the text sample is fully LLM-generated. For this case, we do not consider $y_{i}$ since LLM-generated text is considered as non-LLM-collaboration. We can imply that the similarity between two texts written by the same LLM is higher than that of two LLM families. Hence:
@@
The reverse condition is also true. We can conclude that:
@@
For cases where all samples are human–LLM collaborative, two texts created by the same LLM family tend to be more similar than the ones that involve contributions from different families. That is:
@@
[…]
To enforce the similarity constraints outlined in Eq (6), we build upon the SimCLR framework Chen et al. (2020) and introduce a strategy for defining both positive and negative sample pairs, which forms the basis of our contrastive learning loss. Departing from traditional contrastive losses that rely on a single positive sample, our approach considers a group of positive instances that satisfy specific criteria. The similarity between the anchor and the positive samples is computed as the average similarity across this entire positive set. For negative samples, we follow the methodology used in SimCLR. The resulting contrastive loss, expressed in Eq (7), involves $q$ as the anchor sample, $K^{+}$ as the positive sample set, $K^{-}$ as the negative sample set, $\tau$ as the temperature parameter, and $N_{K^{+}}$ as the number of positive samples.
@@
Therefore, due to the last inequality in Eq (6) only specifying a case ($y_{i,j,k}=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$. Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$. We set $\gamma=\delta=1$, then $\zeta=2,\alpha=\beta=2$. This encourages the model to capture subtle and detailed features from different sources. As a result, it becomes more adept at recognizing variations in writing styles. This capability enhances its accuracy and strengthens its ability to generalize when detecting LLM-generated text.
@@ 4.3 Multi-Task Auxiliary Learning
Multi-task learning Caruana (1997) allows a model to learn several tasks concurrently by sharing relevant information across them. This joint learning process helps the model develop more general and distinctive features. Therefore, it improves the model’s generalizability to new data. Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer. This classifier is responsible for binary classification, determining whether a given text was written by a human or a LLM. Let the probability of $i^{th}$ sample with label $x_{i}=0$ be $p_{i}$. To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, and defined as follows:
@@
Figure 2: Inference architecture: (a) embed the input text into embedding vector using the fine-tuned encoder, (b) use Fuzzy kNN to cluster, retrieving which cluster the input text belongs to (see more in Appendix F), (c) the stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder. If the input text is unseen, we embed it and save it into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
@@ 4.4 Handling Unseen Data without Retraining
Unseen data, whether from an unseen domain or an unfamiliar generator, remains a significant challenge even for state-of-the-art LLM-generated text detection methods, due to the increasing capabilities of LLMs. We tried to use the model classifier on its own, but we ended up using a vector database along with Fuzzy k-Nearest Neighbors, as illustrated in Figure 2. The results were written in Appendix F. Specifically, when dealing with unseen data, we use our model to embed these samples and add them to our existing vector database. Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without the need retraining.
@@ 5 Experiments
In this section, we describe the datasets and baselines we used, followed by two experiments evaluating FAID: (i) classify a text as human, LLM, and human–LLM, and (ii) identify specific generators.
@@
In-domain,
Known generator
FAIDSet LLM-DetectAIve 94.34 94.45 93.79 94.10 0.1888 0.1107
@@
Unseen domain,
Unseen generator
Unseen domain LLM-DetectAIve 52.83 47.31 64.62 53.28 0.4733 0.4722
@@
Unseen
generators
LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
@@
Unseen domain
and Unseen
generators
LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
@@
Table 1: Performance with three labels. For in-domain and known generator datasets, we train the detectors on the training set and evaluate them on the test set. For unseen data evaluation, we use the detector trained on FAIDSet and evaluate on the unseen datasets. The best results are in bold and the second best are underlined.
@@ 5.1 Datasets
In addition to FAIDSet, for in-domain evaluation, we used two additional datasets:
LLM-DetectAIve Abassy et al. (2024) encompasses various domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews. The original human-written and machine-generated texts were augmented by multiple LLMs to create a 236k-example dataset with two new labels: (imachine-written then machine-humanized, and (iihuman-written then machine-polished.
HART Bao et al. (2025) with 21,500 examples including student essays, arXiv abstracts, story writing, and news articles. It contains four categories: human-written, AI-refined, AI-generated, and humanized AI-generated texts. The authors further expanded the dataset by creating additional instances with unbalanced label distributions.
To evaluate the generalizability of FAID on unseen scenarios, we collected the following data:
Unseen domain: We created a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written. We used these essays to generate human–LLM collaborative and LLM-generated texts with the same models with FAIDSet.
Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for the remaining labels using three new LLM families: Qwen, Mistral, and Gemma.
Unseen domain & generators: Based on the human-written IELTS essays above, we used the same LLM families as for the unseen generator test set to generate data for the LLM and the human–LLM labels.
@@ 5.2 Baselines
We adapted the method of Abassy et al. (2024) by fine-tuning a roberta-base sequence classification model. The input texts were tokenized using the RoBERTa tokenizer with a maximum sequence length of 256, and the model was trained for three-class detection.
@@
We adapt the T5-Sentinel framework introduced by Chen et al. (2023) for our three-class setup. Following the original configuration, we train using the AdamW optimizer (batch size 128, learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-5}$). This allows direct comparison with prior T5-based detectors under our experimental conditions.
SeqXGPT:
Following Wang et al. (2023) by modeling token-level likelihood patterns from LLM tokenizers for sentence-level detection, we adapted their method by updating the tokenizer models to align with our dataset’s label space. This adjustment ensures that the extracted token-level log-probability features correspond more accurately to the model types represented in our data.
5.3 Human, LLM or Human-LLM?
Table 1 shows the performance of FAID and three baselines in three evaluation settings.
FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domains, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii). Although SeqXGPT aimed at extracting sequence-level features, it struggles with texts from advanced models, whose coherent and human-like writing styles reduce detectable distinctions.
FAID improves generalization performance over unseen domains and generators compared to others. We can see that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55%, respectively. These results suggest that FAID is an effective method for addressing the multilingual fine-grained LLM-generated text detection task, improving performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to LLM families, rather than overfitting to surface-level artifacts.
@@ 5.4 Identifying Different Generators
We further examine the detector to identify specific AI generators seen during training. The goal of FAID is not only to detect AI involvement, but also to identify the specific LLM or family applied, treating them as distinct authors. As shown in Table 2, across the three datasets, FAID consistently achieves higher performance compared to other baselines in almost all metrics. LLM-DetectAIve achieves comparable results with our detector FAID on the test set of LLM-DetectAIve dataset, except that its precision is slightly lower. FAID’s high performance when dealing with text originating from diverse known generators within these datasets indicates that FAID learned unique writing patterns and features of different AI generators by leveraging multi-level contrastive learning.
6 Conclusion and Future Work
We presented FAIDSet, a multilingual fine-grained LLM-generated text detection dataset comprising 84k examples, and FAID, a framework designed to distinguish between human-written, LLM-generated, and human–LLM collaborative texts. FAID integrates multi-level contrastive learning and multi-task auxiliary objectives, treating LLM families as stylistic “authors,” which enables it to capture subtle linguistic and stylistic cues that generalize effectively across domains and evolving generative systems. Moreover, its Fuzzy KNN-based inference and training-free incremental adaptation contribute to strong robustness and adaptability to unseen data.
Our experiments demonstrated that FAID consistently outperforms competitive baselines across multiple datasets and settings. Its ability to detect nuanced collaborative writing and to adapt to emerging generative models highlights its potential for real-world deployment.
In the future, we plan to extend FAIDSet to cover more languages, generators, and domains, particularly informal genres such as social media and student writing, to enhance cross-lingual and domain robustness further. We also aim to incorporate adversarial LLM-generated texts and more complex forms of human–LLM collaboration to capture the evolving dynamics of AI-assisted text creation.
@@ 7 Limitations
While FAID demonstrates strong performance and generalization across various domains and LLMs, several limitations remain. First, although our dataset is multilingual and multi-domain, it is still limited in terms of low-resource languages and niche writing domains, which may affect performance in those contexts. FAIDSet is synthetic by construction. The controlled generation enables causal-style analysis, but it under-represents the messy, tool-chain-specific edits, wich are seen in the wild. We mitigate this with diverse prompt paraphrases, manual spot-checks, and generalization tests to held-out generators. Second, our framework is based on the observation that texts produced by LLMs from the same family share similar stylistic features. However, this may break down when a single text is influenced by multiple LLM systems – such as when a human uses different models for drafting, rewriting, and polishing a text. In such cases, the resulting style may blend traits from multiple LLMs, making it more difficult to attribute authorship to a single LLM family or clearly distinguish collaboration boundaries.
8 Ethics and Broader Impact
@@
A primary ethical consideration is the data license. We reused existing datasets for our research: LLM-DetectAIve, HART, and IELTS Writing, which have been publicly released under clear licensing agreements. We adhere to the intended usage of these dataset licenses.
@@
FAIDSet streamlines both the creation and the rigorous testing of FAID. By spotting LLM-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors. More broadly, it supports efforts to prevent the misuse of generative technologies in areas such as credential falsification. Detecting LLM-generated content across different languages can be tricky, due to language’s grammar and style. By enabling robust, multilingual and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly. At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and limits of generative AI.
@@
AI-generated text detection systems may inadvertently encode or amplify biases present in the training data. FAIDSet has been carefully constructed to include diverse domains and languages to reduce such biases. Nonetheless, we encourage ongoing auditing and benchmarking of fairness across populations and writing styles, and welcome community feedback for further improvements.
@@
We promote open research and community contributions, and thus we publish our code and data.